# Diffusion-Based Electrocardiography Noise Quantification via Anomaly Detection

### v1

@@ Abstract
Electrocardiography (ECG) signals are often degraded by noise, which complicates diagnosis in clinical and wearable settings. This study proposes a diffusion-based framework for ECG noise quantification via reconstruction-based anomaly detection, addressing annotation inconsistencies and the limited generalizability of conventional methods. We introduce a distributional evaluation using the Wasserstein-1 distance ($W_{1}$), comparing the reconstruction error distributions between clean and noisy ECGs to mitigate inconsistent annotations. Our final model achieved robust noise quantification using only three reverse diffusion steps. The model recorded a macro-average $W_{1}$ score of 1.308 across the benchmarks, outperforming the next-best method by over 48%. External validations demonstrated strong generalizability, supporting the exclusion of low-quality segments to enhance diagnostic accuracy and enable timely clinical responses to signal degradation. The proposed method enhances clinical decision-making, diagnostic accuracy, and real-time ECG monitoring capabilities, supporting future advancements in clinical and wearable ECG applications.
@@ Introduction
Although various approaches have been proposed to address ECG noise, the current methodologies exhibit notable limitations. Traditional denoising methods such as Kalman filtering [10], wavelet transforms [11], and empirical mode decomposition [12] rely on fixed assumptions about the signal or noise characteristics. More recently, deep learning-based denoising methods have been trained to learn noise removal by adding synthetic noise to clean ECG signals [13, 14, 15]. However, these methods assume the presence of clean reference signals, which limits their performance when the ECGs are severely corrupted or missing, resulting in unrealistic reconstructions. Classification-based approaches, including rule-based algorithms [16, 17], signal quality indices (SQIs) [18, 19, 20, 21], and deep feature extractors [22, 23], are also constrained by annotation inconsistencies and label noise. These limitations underscore the need for more scalable and generalizable frameworks specifically designed for ECG noise quantification in realistic settings.
Given the unpredictable nature of noise and artifacts in ECG signals, their detection can be framed as an anomaly detection task. Among the various anomaly detection strategies [24, 25, 26, 27], reconstruction-based approaches have garnered considerable interest owing to their interpretability, which allows for direct comparison between reconstructed and original samples while also modeling in-distribution data without requiring explicit labels or prior knowledge of anomalies [28, 29, 30]. In these methods, a model is trained to reconstruct in-distribution samples, flagging inputs that yield high construction errors during inference as potential anomalies. However, these methods have limitations related to the information bottleneck. If the bottleneck is too narrow, in-distribution samples cannot be properly reconstructed; if it is too wide, out-of-distribution inputs may also be reconstructed with a low error, thereby blurring the distinction between normal and abnormal data [31].

### v2

@@ Abstract
Electrocardiography (ECG) signals are frequently degraded by noise, limiting their clinical reliability in both conventional and wearable settings. Existing methods for addressing ECG noise, relying on artifact classification or denoising, are constrained by annotation inconsistencies and poor generalizability. Here, we address these limitations by reframing ECG noise quantification as an anomaly detection task. We propose a diffusion-based framework trained to model the normative distribution of clean ECG signals, identifying deviations as noise without requiring explicit artifact labels. To robustly evaluate performance and mitigate label inconsistencies, we introduce a distribution-based metric using the Wasserstein-1 distance ($W_{1}$). Our model achieved a macro-average $W_{1}$ score of 1.308, outperforming the next-best method by over 48%. External validation confirmed strong generalizability, facilitating the exclusion of noisy segments to improve diagnostic accuracy and support timely clinical intervention. This approach enhances real-time ECG monitoring and broadens ECG applicability in digital health technologies.
@@ Introduction
Although various approaches have been proposed to address ECG noise, the current methodologies exhibit notable limitations. Traditional denoising methods such as Kalman filtering [10], wavelet transforms [11], and empirical mode decomposition [12] rely on fixed assumptions about the signal or noise characteristics. More recently, deep learning-based denoising methods have been trained to learn noise removal by adding synthetic noise to clean ECG signals [13, 14, 15]. However, these methods assume the presence of clean reference signals, which limits their performance when the ECGs are severely corrupted or missing, resulting in unrealistic reconstructions. Classification-based approaches, including rule-based algorithms [16, 17], signal quality indices (SQIs) [18, 19, 20, 21], and deep feature extractors [22, 23], are also constrained by annotation inconsistencies and label noise. These challenges highlight the need for a methodological shift. Here, we reframe ECG noise quantification as an anomaly detection problem. Instead of training a model to recognize a finite set of predefined noise categories, we train it to capture a comprehensive representation of clean ECG morphology. Any signal that deviates from this learned normative distribution is consequently identified as noise without requiring explicit artifact labels. This approach is intrinsically more robust to the unmodeled and unpredictable artifacts encountered in real-world clinical and ambulatory settings.
Among the various anomaly detection strategies [24, 25, 26, 27], reconstruction-based approaches have garnered considerable interest owing to their interpretability, which allows for direct comparison between reconstructed and original samples while also modeling in-distribution data without requiring explicit labels or prior knowledge of anomalies [28, 29, 30]. In these methods, a model is trained to reconstruct in-distribution samples, flagging inputs that yield high construction errors during inference as potential anomalies. However, these methods have limitations related to the information bottleneck. If the bottleneck is too narrow, in-distribution samples cannot be properly reconstructed; if it is too wide, out-of-distribution inputs may also be reconstructed with a low error, thereby blurring the distinction between normal and abnormal data [31].